OpenStudy (anonymous):

I have some parts of a triangle but I don't know how to use them, can I have help piecing it together?

OpenStudy (anonymous):

Parts as in the lengths? or some lengths and some angles? or what?

OpenStudy (anonymous):

I need to find sin theta, when cot theta is -2 and cos theta is < 0

hartnn (hartnn):

using triangle only ?? you can use identities , if u know them

OpenStudy (anonymous):

hlaf angleS?

hartnn (hartnn):

no, half angle identities won't be required first u find cos theta, using pythgorean identity, then cot theta = cos theta/sin theta

hartnn (hartnn):

oh, sorry

OpenStudy (anonymous):

ok I'll try it

hartnn (hartnn):

first u find cosec theta, using pythgorean identity, then sin theta =1/ cosec theta

OpenStudy (anonymous):

Proceed with the triangle. since each trig function is a ratio of sides on a triangle you can draw what you know, complete the final side and find your last ratio. The only part that requires thought is what quadrant it will be in (based on the signs) 

OpenStudy (anonymous):

so 2(cot^2 theta)=1/cos^2 theta?

OpenStudy (anonymous):

 \[\cot\theta=\frac{adj}{opp}\] {fill in first triangle, solve the hypotenuse} \[\sin\theta=\frac{opp}{hyp}=\frac{1}{\sqrt5}\]